OpenStudy (anonymous):

Parrallelogram perimeter=62 m. base=10.5 m. slant side=____ help me please?

OpenStudy (anonymous):

each slant side is 20.5

OpenStudy (lgbasallote):

perimeter = 2(b + slant height) 62 = 2b + 2slant height 62 = 21 + 2 slant height 41 = 2 slant height 20.5 = slant height

OpenStudy (anonymous):

(10.5*2)=21 62-21=41 41/2=20.5